A +200μC is place at the origin and a -300μC is placed on 1m to the right of it. What is the magnitude and direction of the electric field midway between the 2 charges and 30 cm to the right of the negative charge?

1
Expert's answer
2022-03-08T08:56:38-0500


The net field

"E=E_1+E_2=k\frac{|q_1|}{r_1^2}+k\frac{|q_2|}{r_2^2}"

"E=9*10^9\left(\frac{200*10^{-6}}{0.5^2}+\frac{300*10^{-6}}{0.5^2}\right)\\
=1.8*10^7\:\rm N/C"